# Multilingual Performance Biases of Large Language Models in Education

### v1

@@ 4 Related Work
LLMs, trained on vast multilingual datasets, have demonstrated strong performance in tasks such as text generation, translation, and dialogue Brown et al. (2020), making them promising tools in Intelligent Tutoring Systems (ITS) (Corbett et al., 1997; Pal Chowdhury et al., 2024). Prior work explores their use in educational contexts, such as dynamic student interactions Schmucker et al. (2023), simulating expert and novice behavior Liu et al. (2023), and reasoning in math word problems Opedal et al. (2023).
Beyond mathematical context, LLMs have also been explored for other forms of learning. Cui and Sachan (2023) investigates LLMs in adaptive and personalized exercise generation for language learners, while Wang et al. (2023) examines how conversational tutoring strategies can aid student understanding. Additionally, LLMs have been used to assess grammatical correctness and translation accuracy Kocmi and Federmann (2023); Omelianchuk et al. (2024); Freitag et al. (2024), facilitate automated essay scoring Pack et al. (2024), and provide corrective feedback in second language writing Han et al. (2024).
@@
Recent work has begun addressing multilingual education more directly. Projects like Kaleidoscope Salazar et al. (2025) and Aya Üstün et al. (2024) by Cohere For AI aim to support culturally diverse languages, while SEA-HELM Susanto et al. (2025) and ECLeKTic Goldman et al. (2025) emphasise culturally grounded evaluations in Southeast Asian and cross-lingual contexts, respectively. These efforts highlight the need for multilingual benchmarks that move beyond English-centric evaluations.
@@ 5 Conclusion
We analyse the performance of six well-known state-of-the-art LLMs across six languages other than English on four educational tasks. We find that while performance in English continues to be better than other languages, the drop to other models is not always Large. Particularly, we find that GPT4o and Gemini 2.0 perform consistently well across all languages with a few exceptions. We also note that English prompts work as well, if not better, than prompts written in the target language when solving multilingual tasks. This opens up opportunities for easily porting applications developed for English into different languages. However, we note that certain models can do terribly on some tasks and languages, so we recommend first verifying that a particular model works well in a particular language on a specific education-related task before deployment.
@@
The shown experiments could naturally be better extended to more languages. The selected languages reflect a balance between author familiarity, which is necessary for meaningful qualitative analysis, and linguistic diversity, as evidenced by their spread in URIEL feature space. Similarly, we only covered six LLMs. In both cases, the cost of experiments (see Table 7) becomes prohibitively expensive, which motivated the data release in this paper to enable further research.
@@ A.4 Task: Translation Grading
This sequence is repeated twice for each sentence, once with the original translation and once with the perturbed translation. The scores are then compared. When using English prompts, the LANGUAGE fields are set to their English exonyms, i.e., Hindi, Arabic, Farsi, Telugu, Ukrainian, and Czech. When using translated prompts, the System Prompt, User Message 2, and Gemini Message are translated to the target language. We manually made sure that the formatting was maintained after the translation. We also use the language endonyms, namely [Uncaptioned image].
@@ Appendix B Translation Quality
In the case of Persian, the only recurring error was with mathematical notation, particularly that the minus sign gets placed to the right of the numbers instead of the left, where it should be. This, however, seems to be a rendering issue, which is a result of the fact that the minus sign (‘--’, U+2212) is often replaced by the similar-looking hyphen (‘-’, U+002D), confusing the rendering program into believing that it is rendering text. This should not be an issue since LLMs take raw Unicode encodings as input. Beyond this, there were some minor tense errors, but the meanings were clear.
@@
Figure 6: Evaluation results of the four tasks across five lare language models. The error bars show a 95% confidence interval (t-test). MathDial Graphs show tutoring score after five turns, most models flatline after 5 utterance pairs. The English language column is absent because translation evaluation uses English as the source. All scores range from 0.0 to 1.0, with higher being better, though they are not comparable with each other. Note the truncated y-axes for better detail. Visualizes Tables 6, 6, 6 and 6.
@@
Figure 7: Evaluation results of the four tasks across five lare language models. The error bars show 95% confidence interval (t-test). MathDial Graphs show tutoring score after five turns, most models flatline after 5 utterance pairs. The English language column is absent because translation evaluation uses English as the source. All scores range from 0.0 to 1.0, with higher being better, though they are not comparable with each other. Note the truncated y-axes for better detail. Visualizes Tables 6, 6, 6 and 6.

### v2

@@ 4 Related Work
LLMs, trained on vast multilingual texts, have dominated tasks such as text generation, translation, and dialogue Brown et al. (2020), making them promising tools in Intelligent Tutoring Systems (ITS; Corbett et al., 1997; Pal Chowdhury et al., 2024). Prior work explores their use in educational contexts, such as dynamic student interactions Schmucker et al. (2023), simulating expert and novice behavior Liu et al. (2023), and math word problem reasoning Opedal et al. (2023).
Beyond mathematical context, LLMs have also been explored for other forms of learning. Cui and Sachan (2023) investigate LLMs in adaptive and personalized exercise generation for language learners, while Wang et al. (2023) examines how conversational tutoring strategies can aid student understanding. Additionally, LLMs have been used to assess grammatical correctness and translation accuracy Kocmi and Federmann (2023); Omelianchuk et al. (2024); Freitag et al. (2024), facilitate automated essay scoring Pack et al. (2024), and provide corrective feedback in second language writing Han et al. (2024).
Model API Total Miconception Feedback Tutoring Translation
Mistral Mistral API \$530 \$170 \$170 \$120 \$70
Claude Anthropic \$600 \$190 \$190 \$135 \$85
Command Cohere \$520 \$165 \$165 \$120 \$70
Llama Together.ai \$600 \$190 \$190 \$135 \$80
GPT4o Open AI \$80 \$25 \$25 \$18 \$12
Gemini Google Genai \$30 \$10 \$10 \$6 \$4
Table 8: Approximate costs for the experiments. Does not include taxes or currency conversion charges. The total is about \$2360 with approximately an additional \$500 spent on preliminary experiments.
@@
Recent research has increasingly explored the application of LLMs in multilingual educational contexts, though challenges persist in balancing performance across languages. Systematic reviews of AI-based language learning tools highlight the prevalence of NLP and machine learning techniques for error correction, feedback provision, and assessment in non-English contexts, though they note persistent gaps in dialogic competence and teacher preparedness Alhusaiyan (2025). Studies evaluating LLMs’ cross-lingual capabilities reveal performance disparities, with models demonstrating stronger skill tagging accuracy for English-centric curricula compared to underrepresented languages like Irish or Marathi Kwak and Pardos (2024). Bibliometric analyses indicate growing research interest in AI for foreign language education, particularly in vocabulary acquisition and writing support, though most studies still focus on high-resource European and Asian languages Doğan and Talan (2024). These works collectively underscore both the transformative potential and current limitations of LLMs in achieving equitable multilingual educational support.
To address multilingual education more directly, projects like Kaleidoscope Salazar et al. (2025) and Aya Üstün et al. (2024) by Cohere For AI aim to support culturally diverse languages, while SEA-HELM Susanto et al. (2025) and ECLeKTic Goldman et al. (2025) emphasise culturally grounded evaluations in Southeast Asian and cross-lingual contexts, respectively. These efforts highlight the need for multilingual benchmarks that move beyond English-centric evaluations.
@@ 5 Conclusion
We analyse the performance of six well-known state-of-the-art LLMs across six languages other than English on four educational tasks. We find that while performance in English continues to be better than in other languages, the drop to other models is not always large. In particular, we find that GPT4o and Gemini 2.0 perform consistently well across all languages, with a few exceptions. We also note that English prompts work as well, if not better, than prompts written in the target language, when solving multilingual tasks. This opens up opportunities for porting applications developed for English into different languages. However, we note that certain models perform poorly in some tasks and languages, so we recommend first verifying that a model works well in a particular language on a specific educational task before deployment. However, to answer the question posed by the title, we believe that atleast some language models are reliable across languages.
@@
The shown experiments could naturally be better extended to more languages. The selected languages reflect a balance between author familiarity, which is necessary for meaningful qualitative analysis, and linguistic diversity, as evidenced by their spread in URIEL feature space. Similarly, we only covered six LLMs. In both cases, the cost of experiments (see Table˜8) becomes prohibitively expensive, which motivated the data release in this paper to enable further research.
@@
Acknowledgements
Sankalan Pal Chowdhury is partially funded by the ETH-EPFL JDPLS Program. Donya Rooein is supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 949944, INTEGRATOR).
@@ A.4 Task: Translation Grading
This sequence is repeated twice for each sentence, once with the original translation and once with the perturbed translation. The scores are then compared. When using English prompts, the LANGUAGE fields are set to their English exonyms, i.e., Mandarin, Hindi, Arabic, German, Farsi, Telugu, Ukrainian, and Czech. When using translated prompts, the System Prompt, User Message 2, and Gemini Message are translated to the target language. We manually made sure that the formatting was maintained after the translation. We also use the language endonyms, namely [Uncaptioned image].
@@ Appendix B Translation Quality
In the case of Persian, the only recurring error was with mathematical notation, particularly that the minus sign gets placed to the right of the numbers instead of the left, where it should be. This, however, seems to be a rendering issue, which is a result of the fact that the minus sign (‘-’, U+2212) is often replaced by the similar-looking hyphen (‘-’, U+002D), confusing the rendering program into believing that it is rendering text. This should not be an issue since LLMs take raw Unicode encodings as input. Beyond this, there were some minor tense errors, but the meanings were clear.
@@
Figure 6: Evaluation results of the four tasks across five lare language models. The error bars show a 95% confidence interval (t-test). MathDial Graphs show tutoring score after five turns, most models flatline after 5 utterance pairs. The English language column is absent because translation evaluation uses English as the source. All scores range from 0.0 to 1.0, with higher being better, though they are not comparable with each other. Note the truncated y-axes for better detail. Visualizes Tables˜6, 6, 6 and 6.
@@
Figure 7: Evaluation results of the four tasks across five large language models. The error bars show 95% confidence interval (t-test). MathDial Graphs show tutoring score after five turns, most models flatline after 5 utterance pairs. The English language column is absent because translation evaluation uses English as the source. All scores range from 0.0 to 1.0, with higher being better, though they are not comparable with each other. Note the truncated y-axes for better detail. Visualizes Tables˜6, 6, 6 and 6.